Show that the quantity of labor(X1) and capital(X2) that a firm demands decreases with a factor’s own factor price (w for labor and r for capital) and increases with the output price (P) when the production function is a Cobb-Douglas of the form q=AX1αX2β


1
Expert's answer
2021-10-17T16:50:39-0400

"q=AX_1^\alpha X_2^\beta"

"\alpha" denotes the elasticity of production of labor.

"\beta" denotes the elasticity of production of capital.

A is an implication of the technology used in the production process. When the value of A is high, the level of output that can be produced by any combination of the outputs is also high.

The Cobb Douglas function given is a homogeneous function and the degree of homogeneity is given by: "\alpha" +"\beta" .

"A(tX_1)^\alpha (tX_2)^\beta=t^{\alpha-\beta}AX_1^\alpha X_2^\beta"

"=t^{\alpha-\beta}".

t is a real positive value.

From the above equation, we get the implication that if L and K by the factor t, Q will increase by the factor "t^{\alpha-\beta}"

In increase in factor prices of labor and capital will result in less of them being demanded by the firm.

However, when Q increases by the factor "t^{\alpha-\beta}",the firm's demand for the input factors also increases by "t^{\alpha-\beta}" .